# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v2

@@ 1 Introduction
Modern generative AI models, pre-trained on extensive multi-modal datasets, can create synthetic images (e.g., Stable Diffusion [65], DALL-E [64], MidJourney [52]), videos (e.g., OpenAI Sora [60], Microsoft VASA-1 [81]), audio (e.g., AudioGen [31]) and text/conversations (e.g., ChatGPT [5], BlenderBot [67]). These models have enabled a range of new applications such as creative content generation, personalized virtual assistants, and sophisticated chatbots. Foundational Large Language Models (LLMs), including GPT-4 [1] and LlaMA [74], are trained on enormous and diverse textual datasets, and typically excel in a variety of natural language tasks, such as answering questions, summarizing documents, and translating languages.
@@
An emerging application of these LLMs is to generate computer code, which is typically accomplished by first training or fine-tuning them using vast amounts of programming/code datasets found on online repositories (e.g., GitHub), technical forums, and documentation. Both commercial/black-box (e.g., GPT-4 [1], Claude [3]) and open-source (e.g., CodeLlama [66], DeepSeek Coder [17]) varieties of such code-generating LLMs are readily available and being extensively used by both novice and expert programmers in their coding workflows to increase productivity. Recent studies indicate that up to 97% of developers are using generative AI to some degree and that approximately 30% of code written today is AI-generated, reflecting significant perceived gains in efficiency and convenience [68, 41].
@@
One serious shortcoming of LLMs is a phenomenon referred to as hallucination. Hallucinations are outputs produced by LLMs that are factually incorrect, nonsensical, or completely unrelated to the input task. Hallucinations present a critical obstacle to the effective and safe deployment of LLMs in public-facing applications due to their potential to generate inaccurate or misleading information. As a result, there has been increased efforts to research the detection and mitigation of hallucinations in LLMs [20, 26]. However, most existing research has focused only on hallucinations in classical natural language generation and prediction tasks such as machine translation, summarization, and conversational AI [22, 51, 38, 9]. The occurrence and impact of hallucinations during code generation, particularly regarding the type of hallucinated content and its implications for code security, are still in the nascent stages of research. Recently, Liu et al. [44] have shown that popular LLMs (e.g., ChatGPT, CodeRL, and CodeGen) significantly hallucinate during code generation and have established a taxonomy of hallucinations in LLM-generated code.
@@
In this work, we focus on a specific type of hallucination during code generation called package hallucination. Package hallucination occurs when an LLM generates code that recommends or contains a reference to a package that does not actually exist. An adversary can exploit package hallucinations, especially if they are repeated, by publishing a package to an open-source repository with the same name as the hallucinated or fictitious package and containing some malicious code/functionality. Then, as other unsuspecting and trusting LLM users are subsequently recommended the same fictitious package in their generated code, they end up downloading the adversary-created malicious package, resulting in a successful compromise. This compromise can then spread through an entire codebase or software dependency chain, infecting any code that relies on the malicious package. This is a variation of the classical package confusion attack that has been enabled by code-generating LLMs.
@@
Package confusion attacks, through techniques such as typosquatting (i.e., creating packages with names similar to popular ones to deceive users) and name similarity, have been a longstanding issue in the open-source software community [55, 69, 32]. Package hallucinations by code-generating LLMs threaten to exacerbate the problem by exposing an additional threat surface for such attacks. Trivial cross-referencing methods (i.e., comparing a generated package name with a list of known packages) are ineffective for detecting a package hallucination attack, as an adversary may have already published the hallucinated package with malicious code. Open source repositories make no guarantee about the safety of hosted content; the mere presence of a package in an open-source repository does not confirm its credibility. A recent blog post [34] suggests that LLMs are prone to package hallucinations and provides a first approximation of its prevalence, but the extent to which this phenomenon occurs in state-of-the-art (SOTA) commercial and open-source LLMs, the nature of these hallucinations, and the effectiveness of potential mitigation measures have not been comprehensively investigated before.
@@
In this paper we conduct the first systematic study of the frequency and nature of package hallucinations across a variety of code generation LLMs operating under a diverse set of model settings and parameters. We specifically make the following novel contributions:
@@
  • Quantifying the frequency and prevalence of package hallucinations by code generating LLMs: We first comprehensively analyze the prevalence of package hallucinations in Python and JavaScript code generated by popular publicly-available commercial and open-source LLMs. We also examine the common behaviors in LLMs that lead to package hallucinations, including hallucination repetition, output verbosity, and the ability of models to detect their own hallucinations.
@@
  • Analyzing the effect of model settings and training data on package hallucinations: We further study how specific model settings, such as training data recency, model temperature, and decoding strategies affect the occurrence and nature of package hallucinations.
@@
  • Characterizing the common traits of the generated hallucinated packages: We carefully study several key properties of the hallucinated packages, such as their semantic similarity to popular packages, the occurrence of package recommendations from other programming languages, package persistence, and the influence of recently removed packages on hallucination rate.
@@
  • Mitigation Strategies: We propose and comprehensively evaluate several techniques to effectively mitigate package hallucinations in LLM-generated code while maintaining the ability to produce effective code, something that has not been done before.
  • Data We make publicly available a novel dataset of 19,500 coding prompts for a vast range of coding tasks, 586,000 generated coding samples for fine-tuning/analysis, and a list of 205,474 hallucinated package names that can be used to monitor for potential instances of a package hallucination attack.
@@ 2 Background and Related Work
In this section we provide a brief background on open-source software security, code generating LLMs, and the issue of hallucinations in LLMs.
@@ 2.1 Background
Package Confusion Attacks in Open-source Software Repositories. Modern software development has seen an increased reliance on open-source software packages and libraries available on public centralized repositories. Many modern programming languages now rely on such centralized package repositories, with PyPI [12] (for Python) and npm [56] (for JavaScript) being the two most popular repositories. The open nature of these repositories, where anyone can upload new code packages/libraries, makes them an attractive platform for malware distribution. For instance, a total of 245,000 malicious code packages were discovered on open-source software repositories in 2023 alone [68].
@@
Packages/libraries often rely on other packages to function, thus creating extensive dependency trees, infecting a single package in this dependency chain can be sufficient to compromise an entire software product or ecosystem [55, 30]. Public OSS repositories such as PyPI and npm have implemented various measures, including two-factor authentication, namespace protection, and software signing to mitigate the distribution of malicious packages [76, 82]. However, it remains unclear whether these repositories utilize any scan-based techniques for detecting malicious code, and they often do not disclose the full list of removed packages.
Automated Code Generation using LLMs. Modern LLMs are continuing to demonstrate advanced source code generation capabilities, with success rates for correctly answering coding prompts surging from 25% in June 2021 to 96% by April 2024 [8]. With the increasing use of these models for software development, concerns are rising about the likelihood of producing insecure or incorrect code that could create vulnerabilities in deployed applications. Early versions of code generation LLMs were found to generate code containing vulnerabilities listed in the MITRE Top-25 Common Weakness Enumeration (CWE) 40% of the time [61]. Moreover, recent research has shown that AI-assisted programming not only results in less secure code but also instills a false sense of security among developers [62].
Hallucinations by LLMs. It has been well-documented that LLMs can unintentionally produce harmful information [46, 79], be manipulated for malicious purposes [18, 29], expose private information [39], and carry inherent biases in their training data [14]. A related phenomenon is hallucinations, where LLMs generate misleading or entirely fictitious information. These errors take various forms: the model might misinterpret the intended input (input-conflicting hallucination), produce inconsistencies with prior outputs (context-conflicting hallucination), or contradict established facts (fact-conflicting hallucination) [83]. Hallucinations can arise from three main root causes: (i) data, (ii) training, and (iii) inference [21]. Data-related hallucinations occur when the source data itself is flawed with misinformation [42], bias [14], or incomplete records [58]. Architecture flaws [43] or suboptimal training objectives [77] during training could also result in downstream hallucinations, while inference time issues such as defective coding strategies [19] and imperfect decoding representations [7, 47] are further contributors. The probabilistic nature of LLMs presents a challenge in mitigating hallucinations. This nondeterminism, while it fosters creativity and generates diverse and innovative content, also contributes to the generation of hallucinations. Balancing creativity with accuracy remains a central challenge in deploying LLMs, underscoring the complexity of developing effective mitigation strategies.
@@ 2.2 Related Work
The idea of code generation models recommending malicious or typosquatted packages was first suggested in 2021 as tools such as GPT-3 and Codex were released as viable code generation platforms [5]. At the time, the risk of code generation tools suggesting vulnerable, malicious, or typosquatted packages was assessed to be low [8]. In particular, the related but distinct concept of package hallucinations was not explicitly considered in this initial risk assessment; either because such an attack had not been considered or because the threat was thought to be negligible. The capabilities of generative AI agents have advanced significantly since that introductory evaluation.
Although a comprehensive study on the prevalence of package hallucinations in LLM-generated code has not been done previously, a recent blog post by Lanyado [35] outlines the results of some preliminary tests on commercial LLMs such as GPT, Gemini, and Cohere. Their testing does confirm the presence of hallucinated packages in the code generated by these LLMs but their initial estimate of hallucination rate is 5x higher than our findings and they do not consider popular open-source LLMs or study possible mitigation approaches. They also fail to systematically characterize the generated hallucinated packages and model properties that have an impact on hallucinations.
In contrast, we conduct a rigorous and comprehensive evaluation across a broader range of models, including the first analysis of package hallucinations in open-source models of any kind, at a scale that has not been done previously. To this end, we provide thorough testing with a larger custom dataset covering two programming languages (namely, Python and Javascript), followed by a detailed analysis of the significant characteristics of this phenomenon.
@@ 3 Research Questions
Adversary Model and Assumptions. We assume an adversary that wants to execute a package confusion attack by leveraging package hallucinations in code generated by both closed-source and open-source code generation LLMs (see Figure 1). The target of the adversary are users who employ such LLMs for generating code. Here we assume that the LLM generates code that requires additional packages for the user to install, which the user does without sufficient verification of the recommended packages. In other words, the target users fully trust the LLMs to include only valid package names in the generated code. We assume that the adversary has access to the same set of LLMs for code generation as the target users, and is unable to modify or manipulate the model and model parameters of these LLMs (e.g., via re-training or fine-tuning) before they are used by the victims.
The adversary is able to determine a list of hallucinated packages generated by these LLMs (for example, by cross-referencing the package repository), and then is able to realize a package confusion attack by creating a package of the same name on the corresponding package repositories.
These newly created (and now publicly available) packages by the adversary could contain malicious code or functionality. Research has shown that installing open source Python or JavaScript packages allows arbitrary code execution to an attacker [57]. Previous work [35] has also established the viability of such an attack by publishing a hallucinated package to an open source repository and demonstrating that the package is actively downloaded and was incorporated into the dependency chains of other packages/code.
Research Questions. We now organize our investigation into five broad Research Questions (RQ).
@@
RQ3: What are the commonly observed model behaviors related to package hallucinations? This RQ will exhaustively study model behaviors such as hallucination repetition by a single LLM (hallucination persistence) and across multiple LLMs (cross-model hallucinations), output verbosity, and the ability of LLMs to detect their own hallucinations.
@@
RQ5: Is it possible to effectively mitigate package hallucinations using best practices in the literature and knowledge gained from earlier results? Through this RQ, we will investigate if code generating LLMs can be designed to reduce hallucinations with minimal compromise to code quality. In this direction, we will study if techniques such as retrieval augmented generation (RAG), self-detected feedback, decoding strategies, and supervised fine-tuning are effective hallucination reduction strategies.
@@ 4 Experiment Design
To address the RQs outlined above, we design several experiments to repeatedly prompt LLMs to generate code and then analyze the generated code. Our experimentation pipeline comprises of three distinct phases: (i) prompt dataset generation, (ii) code generation, and (iii) hallucination detection, each of which is described below:
@@ 4.1 Prompt Dataset
The experiments are designed to exhaustively test each LLM through a complete range of coding tasks. Existing datasets of benchmark coding prompts contain only a limited number of prompts (e.g. only 164 prompts for both EvalPlus [45] and HumanEval [8]) and lack diversity. Therefore, we create a new code prompt dataset for our experiment that contains both breadth and depth in terms of overall number of prompts and range of topics. Our goal was to develop a dataset that accurately and comprehensively represents coding tasks commonly requested by everyday users. To accomplish this, we employed two distinct approaches, as described below.
@@
Stack Overflow Dataset. To model the input prompts around real programmer questions, our first prompt dataset was created using Stack Overflow [70] questions across relevant programming topics and subject areas. Stack Overflow is a popular online question-and-answer service for software programmers and developers. To capture a wide range of topics, we utilized the “tag” feature of Stack Overflow, which allows users to label posts according to a subject matter. We included any tag that had over 5,000 questions and was also relevant to Python or JavaScript (the two programming languages we focus in this work, as detailed in Section 4.2). For each of the 240 manually selected tags that met this criteria (full list of tags is listed in Appendix H, we extracted the 20 most up-voted questions, resulting in 4,800 prompts (i.e., 4,800 prompts for Python and 4,800 prompts for JavaScript).
@@
As more recent data is less likely to be included in the pre-training data of LLMs, we are also interested in investigating the correlation between data recency (i.e., how recently the question was asked on Stack Overflow) and model hallucination rate. To enable such an analysis, we ran two queries on Stack Overflow; one that only captured the most popular questions in the selected tags from 2023 and another that captured the most popular questions for all years prior to 2023. By including the two different ranges of time, we effectively doubled the original number of prompts, for a total of 9,600 for each of the two languages.
@@
Not all questions asked on Stack Overflow may involve coding or require code to answer the question. Rather than attempting to filter out such prompts during the code generation phase, which is non-trivial and error-prone, the LLM is asked to answer the question and only provide code if necessary. In the end, this may result in a slightly lesser number of usable LLM-generated code samples, but is more realistic as LLMs are expected to accommodate imperfect user inputs.
@@
LLM-generated Dataset. As a majority of the programming tasks require some library/package, our next idea was to use the package repositories themselves as a good representation of the full spectrum of coding topics. Our goal was to represent as many code generation tasks as possible in one comprehensive dataset. We take the 5,000 most popular Python and JavaScript packages (according to number of downloads) and scrape the official package description as listed on PyPI and npm, respectively. These descriptions are then input individually to the Llama-2 70B model with instructions to generate a coding prompt based on the package description (exact prompt available in Figure 13). This process worked remarkably well and generated roughly 4,800 prompts for Python and JavaScript each, resulting in two datasets of approximately the same size (some packages with no description or descriptions in a non-English language were discarded). Similar to the Stack Overflow dataset, we doubled the LLM-generated dataset for temporal analysis by dividing it into two segments: packages most downloaded in 2023 and packages most downloaded prior to 2023. When a package appears in both sets, we remove the package from the latter set to ensure no overlap. Removing duplicates from the latter dataset guarantees that the remaining packages will be those that have increased in popularity during the last year, capturing the desired signal. A truncated list from the LLM-generated dataset can be found in Appendix A.

### v3

@@ 1 Introduction
Recent advances in generative AI, powered by Large Language Models (LLMs) like GPT-4 [1] and LlaMA [60], have revolutionized AI capabilities across modalities, excelling in a wide range of tasks such as image synthesis, text generation, and natural language understanding. One such application is code generation, which is typically accomplished by first training or fine-tuning an LLM using vast amounts of programming data found on online repositories (e.g., GitHub), technical forums, and documentation. Both commercial/black-box (e.g., GPT-4 [1], Claude [3]) and open-source (e.g., CodeLlama [53], DeepSeek Coder [14]) varieties of such code-generating LLMs are readily available and are extensively used by both novice and expert programmers in their coding workflows to increase productivity. Recent studies indicate that up to 97% of the developers are using generative AI to some degree and that approximately 30% of code written today is AI-generated, reflecting significant perceived gains in efficiency and convenience [54, 36].
@@
One critical shortcoming of LLMs is a phenomenon referred to as hallucination. Hallucinations are outputs produced by LLMs that are factually incorrect, nonsensical, or completely unrelated to the input task. Hallucinations present a significant obstacle to the effective and safe deployment of LLMs in public-facing applications due to their potential to generate inaccurate or misleading information. As a result, there has been increased efforts to research the detection and mitigation of hallucinations in LLMs [17, 22]. However, most existing research has focused only on hallucinations in classical natural language generation and prediction tasks such as machine translation, summarization, and conversational AI [19, 45, 33, 7]. The occurrence and impact of hallucinations during code generation, particularly regarding the type of hallucinated content and its implications for code security, are still in the nascent stages of research. Recently, Liu et al. [39] have shown that popular LLMs (e.g., ChatGPT, CodeRL, and CodeGen) significantly hallucinate during code generation and have established a taxonomy of hallucinations in LLM-generated code.
@@
In this work, we focus on a specific type of hallucination during code generation called package hallucination. Package hallucination occurs when an LLM generates code that recommends or contains a reference to a package that does not actually exist. An adversary can exploit package hallucinations, especially if they are repeated, by publishing a package to an open-source repository with the same name as the hallucinated or fictitious package and containing some malicious code/functionality. As other unsuspecting and trusting LLM users are subsequently recommended the same fictitious package in their generated code, they end up downloading the adversary-created malicious package, resulting in a successful compromise. This compromise can then spread through an entire codebase or software dependency chain, infecting any code that relies on the malicious package. This is a variation of the classical package confusion attack that has been enabled by code-generating LLMs.
@@
Package confusion attacks, through techniques such as typosquatting (i.e., creating packages with names similar to popular ones to deceive users) and name similarity, have been a long-standing issue in the open-source software community [47, 55, 27]. Package hallucinations by code-generating LLMs threaten to exacerbate the problem by exposing an additional threat surface for such attacks. Trivial cross-referencing methods (i.e., comparing a generated package name with a list of known packages) are ineffective for detecting a package hallucination attack, as an adversary may already have published the hallucinated package with malicious code. Open-source repositories make no guarantee about the safety of hosted content; the mere presence of a package in an open-source repository does not confirm its credibility. A recent blog post [29] suggests that LLMs are prone to package hallucinations and provides a first approximation of their prevalence, but the extent to which this phenomenon occurs in state-of-the-art (SOTA) commercial and open-source LLMs, the nature of these hallucinations, and the effectiveness of potential mitigation measures have not been thoroughly investigated before.
@@
In this paper, we conduct the first systematic study of the frequency and nature of package hallucinations across a variety of code-generating LLMs, operating under a diverse set of model settings and parameters. We specifically make the following novel contributions:
@@
  • Characterizing the prevalence of package hallucinations by code-generating LLMs and related functional attributes: We first comprehensively analyze the prevalence of package hallucinations in Python and JavaScript code generated by popular commercial and open-source LLMs. We also examine and characterize commonly observed LLM behaviors related to package hallucinations, including hallucination repetition, output verbosity, and the ability of these models to detect their own hallucinations.
@@
  • Analyzing the effect of fine-grained changes to model settings on package hallucinations: We further study how specific model settings, such as training data recency, model temperature, and decoding strategies affect the occurrence and nature of package hallucinations.
@@
  • Characterizing common traits of the generated hallucinated packages: We carefully study several key properties of the hallucinated packages, such as their semantic similarity to popular packages, their propensity of occurring across different models, and the influence of packages that were recently removed/deleted (from the corresponding repositories) on the hallucination rate, among others.
@@
  • Testing of mitigation strategies: We propose and comprehensively evaluate several techniques to effectively mitigate package hallucinations in LLM-generated code while maintaining the ability to produce effective code.
  • Publicly-accessible datasets for advancing research: We make publicly available two novel datasets (one Python and one JavaScript) of 19,500 coding prompts for a wide range of coding tasks and 586,000 generated coding samples for fine-tuning/analysis.  [^1] All code and datasets can be found at: https://zenodo.org/records/14676377 or https://github.com/Spracks/PackageHallucination
@@ 2 Background and Related Work
In this section, we provide a brief background on open-source software security, code-generating LLMs, and the issue of hallucinations in LLMs.
@@ 2.1 Background
Package Confusion Attacks in Open-source Software Repositories. Modern software development has seen an increased reliance on open-source software packages and libraries that are publicly-available on centralized repositories. Many modern programming languages now rely on such centralized package repositories, with PyPI [9] (for Python) and npm [48] (for JavaScript) being the two most popular repositories. The open nature of these repositories, where anyone can upload new code packages/libraries, makes them an attractive platform for malware distribution. For instance, a total of 245,000 malicious code packages were discovered in open-source software repositories in 2023 alone [54].
@@
Packages/libraries often rely on other packages to function, thus creating extensive dependency trees. Infecting a single package in this dependency chain can be sufficient to compromise an entire software product or ecosystem [47, 26]. Public OSS repositories such as PyPI and npm have implemented various measures, including two-factor authentication, namespace protection, and software signing to mitigate the distribution of malicious packages [61, 66]. However, it remains unclear whether these repositories utilize any scan-based techniques for detecting malicious code, and they often do not disclose the full list of removed packages.
Automated Code Generation using LLMs. Modern LLMs continue to demonstrate advanced source-code generation capabilities, with success rates in correctly answering coding prompts surging from 25% in June 2021 to 96% by April 2024 [6]. With the increasing use of these models for software development, concerns are increasing about the likelihood of producing insecure or incorrect code that could create vulnerabilities in deployed applications. Early versions of code-generating LLMs were found to generate code containing vulnerabilities listed in the MITRE Top-25 Common Weakness Enumeration (CWE) 40% of the time [51]. Moreover, recent research has shown that AI-assisted programming not only results in less secure code, but also instills a false sense of security among developers [52].
Hallucinations by LLMs. It has been well documented that LLMs can unintentionally produce harmful information [41, 64], be manipulated for malicious purposes [15, 25], expose private information [34], and carry inherent biases in their training data [11]. A related phenomenon is hallucinations, where LLMs generate misleading or entirely fictitious information. These errors take various forms: the model might misinterpret the intended input (input-conflicting hallucination), produce inconsistencies with previous output (context-conflicting hallucination), or contradict established facts (fact-conflicting hallucination) [67]. Hallucinations can arise from three main root causes: (i) data, (ii) training, and (iii) inference [18]. Data-related hallucinations occur when the source data itself is flawed with misinformation [37], bias [11], or incomplete records [50]. Architecture flaws [38] or suboptimal training objectives [62] during training could also result in downstream hallucinations, while inference time issues such as defective coding strategies [16] and imperfect decoding representations [5, 42] are other contributors. The probabilistic nature of LLMs presents a challenge in mitigating hallucinations. This nondeterminism, while it fosters creativity and generates diverse and innovative content, also contributes to the generation of hallucinated content. Balancing creativity with accuracy remains a central challenge in deploying LLMs, underscoring the complexity of developing effective mitigation strategies.
@@ 2.2 Related Work
The possibility of code-generating models recommending malicious or typosquatted packages was first suggested in 2021 as tools such as GPT-3 and Codex were released as viable code generation platforms [4]. At the time, the risk of these tools suggesting vulnerable, malicious, or typosquatted packages was assessed to be low [6]. In particular, the related but distinct concept of package hallucinations was not explicitly considered in this initial risk assessment; either because such an attack scenario was not considered at all or because the threat was thought to be negligible. The capabilities of generative AI agents have advanced significantly since that introductory evaluation.
Although a comprehensive study on the prevalence of package hallucinations in LLM-generated code has not been done previously, a recent blog post by Lanyado [30] outlines the results of some preliminary tests on commercial LLMs such as GPT, Gemini, and Cohere. Their testing confirms the presence of hallucinated packages in the code generated by these LLMs, but their initial estimate of hallucination rate was 5$\times$ higher than our findings, and they do not consider popular open-source LLMs or study possible mitigation approaches. They also fail to systematically characterize the generated hallucinated packages and model properties that have an impact on hallucinations.
In contrast, we conduct a rigorous and comprehensive evaluation across a broader range of models, including the first analysis of package hallucinations in open-source models of any kind, at a scale that has not been previously done. To this end, we provide thorough testing with a larger custom dataset covering two programming languages (namely, Python and Javascript), followed by a detailed analysis of the significant characteristics of this phenomenon.
@@ 3 Research Questions
Adversary Model and Assumptions. We assume an adversary who wants to execute a package confusion attack by leveraging package hallucinations in the code generated by closed-source and open-source code-generating LLMs (see Figure 1). The target of the adversary are users who employ such LLMs for generating code. Here we assume that the LLM generates code that requires additional packages for the user to install, which the user does without sufficient verification of the recommended packages. In other words, the target users fully trust the LLMs to include only valid package names in the generated code. We assume that the adversary has access to the same set of LLMs for code generation as the target users, and is unable to modify or manipulate the model and model parameters of these LLMs (e.g., via retraining or fine-tuning) before they are used by the victims. The adversary is able to determine a list of hallucinated packages generated by these LLMs (for example, by cross-referencing the package repository), and then is able to realize a package confusion attack by creating a package of the same name on the corresponding package repositories. These newly created (and now publicly available) packages by the adversary could contain malicious code or functionality. Research has shown that installing open source Python or JavaScript packages allows the execution of arbitrary code by an attacker [49]. Previous work [30] has also established the viability of such an attack by publishing a hallucinated package to an open source repository and demonstrating that the package is actively downloaded and was incorporated into the dependency chains of other packages/code.
Research Questions. We now organize our investigation into the following five broad Research Questions (RQ).
@@
RQ3: What are the commonly observed model behaviors related to package hallucinations? This RQ will exhaustively study model behaviors such as hallucination repetition by a single LLM (hallucination persistence) and across multiple LLMs (cross-model hallucinations), output verbosity, and the ability of LLMs to detect their own hallucinations (after generation).
@@
RQ5: Is it possible to effectively mitigate package hallucinations using best practices in the literature and knowledge gained from earlier results? Through this RQ, we will investigate if code-generating LLMs can be designed to reduce hallucinations with minimal compromise to code quality. In this direction, we will study if techniques such as retrieval augmented generation (RAG) [32], self-detected feedback, decoding strategies, and supervised fine-tuning [57] are effective package hallucination reduction strategies.
@@ 4 Experiment Design
To address the RQs outlined above, we design several experiments to repeatedly prompt LLMs to generate code and then analyze the generated code. Our experimentation pipeline consists of three distinct phases: (i) prompt dataset generation, (ii) code generation, and (iii) hallucination detection, each of which is described next.
@@ 4.1 Prompt Dataset
The experiments are designed to exhaustively test each LLM through a complete range of coding tasks. Existing benchmark datasets of coding prompts contain only a limited number of prompts (e.g. only 164 prompts for both EvalPlus [40] and HumanEval [6]) and lack diversity. Therefore, we create a new code prompt dataset for our experiment that contains both breadth and depth in terms of overall number of prompts and range of topics. Our goal was to develop a dataset that accurately and comprehensively represents the coding tasks commonly requested by everyday users. To accomplish this, we employ two distinct approaches, as described below.
@@
Stack Overflow Dataset. To model the input prompts around real programmer questions, our first prompt dataset was created using Stack Overflow [56] questions across relevant programming topics and subject areas. Stack Overflow is a popular online question-and-answer service for software programmers and developers. To capture a wide range of topics, we utilize the “tag” feature of Stack Overflow, which allows users to label posts according to a subject matter. We included any tag that had more than 5,000 questions and was also relevant to Python or JavaScript (the two programming languages that we focus on in this work, as detailed in Section 4.2). For each of the 240 manually selected tags that met this criterion (a full list of tags can be found with the paper artifacts 1), we extracted the 20 most upvoted questions, resulting in 4,800 prompts (i.e., 4,800 prompts for Python and 4,800 prompts for JavaScript).
@@
As more recent data is less likely to be included in the pre-training data of LLMs, we are also interested in investigating the temporal correlation between data recency (i.e., how recently the question was asked on Stack Overflow) and model hallucination rate. To enable such an analysis, we ran two queries on Stack Overflow; one that captured only the most popular questions in the selected tags from 2023 and another that captured the most popular questions for all years prior to 2023. By including the two different ranges of time, we effectively doubled the original number of prompts, for a total of 9,600 for each of the two languages.
@@
Not all questions asked on Stack Overflow may involve coding or require code to answer the question. Rather than attempting to filter out such prompts during the code generation phase, which is non-trivial and error-prone, the LLM is asked to answer the question and only provide code if necessary. In the end, this may result in a slightly lesser number of usable LLM-generated code samples but is more realistic as LLMs are expected to accommodate imperfect user inputs.
@@
LLM-generated Dataset. As a majority of the programming tasks require some library/package, our next idea was to use the package repositories themselves as a good representation of the full spectrum of coding topics. Our goal was to represent as many code generation tasks as possible in one comprehensive dataset. We take the 5,000 most popular Python and JavaScript packages (based on the number of downloads) and scrape the official package description as listed on PyPI and npm, respectively. These descriptions are then individually inputted to the Llama-2 70B model with instructions to generate a coding prompt based on the package description (the exact prompt available in Figure 13). This process generated roughly 4,800 prompts for Python and JavaScript each, resulting in two datasets of approximately the same size (some packages with no description or descriptions in a non-English language were discarded). Similarly to the Stack Overflow dataset, we doubled the LLM-generated dataset for temporal analysis by dividing it into two segments: the packages most downloaded in 2023 and the packages most downloaded prior to 2023. When a package appears in both sets, we remove the package from the latter set to ensure that there is no overlap. Removing duplicates from the latter dataset guarantees that the remaining packages will be those that have increased in popularity during the last year, capturing the desired signal. A truncated list of the LLM-generated dataset can be found in Appendix A.
For brevity, the two temporally distinct datasets, one from the past year and one from before 2023, will be referred to as the ’recent’ and ’all-time’ datasets for the remainder of this paper.